OpenStudy (anonymous):

Given the following unbalanced chemical equation, how many grams of aluminum chloride are produced when 3.00 moles of aluminum reacts with 5.30 moles of chlorine gas? ___Al + ___ Cl2 arrow ___AlCl3

OpenStudy (anonymous):

471 grams 399 grams 151 grams 177 grams

OpenStudy (chmvijay):

can u tell molecular weight of Al Cl2 and AlCl3

OpenStudy (anonymous):

AI- 26.98153868 CI2- 70.9064 AICI3- 133.3411

OpenStudy (chmvijay):

anc can you balance the reaction and write it

OpenStudy (anonymous):

I do not think so

OpenStudy (chmvijay):

mean u dont know

OpenStudy (anonymous):

yes, I do not know lol

OpenStudy (chmvijay):

:)

OpenStudy (anonymous):

help! lol :)

OpenStudy (chmvijay):

2Al + 3Cl2 = 2AlCl3 is a balanced equation

OpenStudy (anonymous):

wow I was sooooooo off.... how do I figure out grams now?

OpenStudy (chmvijay):

for 2 moles of Al how many moles of Cl2 is required can you tell

OpenStudy (anonymous):

3?

OpenStudy (chmvijay):

yaaa right

OpenStudy (anonymous):

yay! now what?

OpenStudy (chmvijay):

now you have to check which one is limiting reagent

OpenStudy (anonymous):

oh gosh... I would say CI2

OpenStudy (chmvijay):

mean Cl2 is in excess than the Al right bcoz 2 mole of Al needs 3moles of Cl2 from balanced equation but in ur question they have mentioned for 3 moles of Al reacts with 5.30 moles since Cl2 is huge amount hence it is not limiting reagent limiting reagent is Al

OpenStudy (chmvijay):

2mole s of Al give rise too 2 moles of alCl3 3 moles of Al give rise too =3*2 /2 =3 moles of AlCl3 u have to convert moles into grams grams = moles * molecular weight of AlCl3 = 3 * 133.34 = 400 gram

OpenStudy (chmvijay):

can u guess the answer now

OpenStudy (anonymous):

the only one that comes close is 399...

OpenStudy (chmvijay):

yaaa that is the answer LOL:)

OpenStudy (anonymous):

OH!! AWESOME!